OpenStudy (anonymous):

Hard A..Word Problem Plz Help! Members of the Ski Club contributed equally to obtain $1800 for a holiday trip. When 6 members found they could not go, their contributions were refunded and each remaining member then had to pay an additional $10 to remain at their goal of $1800. How many went on the trip?

OpenStudy (anonymous):

answer is 36

OpenStudy (anonymous):

well actually the answer is 30 because 6 dropped out

OpenStudy (anonymous):

work could be something like this: put x = number of original members. then they each paid \[\frac{1800}{x}\] dollars. after 6 drop out they are refunded a total of \[6\times \frac{1800}{x} \] and then the remaining x- 6 pay 10 to make up the balance. you can therefore solve \[1800-6\times \frac{1800}{x}+10(x-6)=1800\]

OpenStudy (anonymous):

or more basically just solve \[10(x-6)=6\times \frac{1800}{x}\] for x